# Designer Spin-Orbit Superlattices: Symmetry-Protected Dirac Cones and Hot Spots of Spin Berry Curvature in 2D van der Waals Metamaterials

### v1

@@ Abstract
The enhanced SB curvature of the SSOF-induced mini-bands indicates that the 2D van der Waals metamaterials proposed here support large spin-Hall responses. To confirm this, we compute the intrinsic spin Hall conductivity ($\sigma_{ij}^{z}$) from the flux of SB curvature using standard methods [69]. According to linear response theory, the $z$-polarized spin current density generated by an external electric field is $\mathbf{j}_{s}^{z}=\sum_{i,j=x,y}\sigma_{ij}^{z}E_{j}\mathbf{e}_{j}$, where $E_{j}$ are the field components and $\textbf{e}_{i}$ is the unit vector along the $i$-axis. As shown in Fig. 3(c), the spin Hall response has a strong energy dependence and can reach sizeable values on the order of $e/4\pi$ for typical values of proximity-induced SOC at room temperature. This strong response is robust to design imperfections and weak disorder provided that the global average of the SSOFs is preserved ($\langle\Phi(x)\rangle=0$). We verified this with different types of proximity-induced SSOFs and spatial patterns $\Phi(x)$. At variance with 2D conductors subject to the usual uniform Rashba effect [72, 49], the spatial dependence of the Rashba SSOF protects our quantum geometry-driven spin Hall effect from exact cancellations due to impurity-scattering corrections. In fact, a SU(2)-covariant conservation law for the spin current can be derived in the vein of Ref. [49] yielding $\langle H_{\textrm{so}}(x)v_{i}^{z}\rangle=0$, with delimited-⟨$\langle$..$\rangle$ … ⟩ denoting a quantum and disorder average. For a uniform Rashba field, this relation (which holds in the presence of arbitrary non-magnetic impurity potentials) implies $\mathbf{j}_{s}^{z}=0$. However, in our system this cancellation is circumvent due to the oscillatory nature of $H_{\textrm{so}}(x)$. The SSOF-driven SHE thus appears to be more robust than its counterpart in standard Rashba-coupled graphene. We now briefly address the case of 2D metamaterials with concurrent Rashba-type and valley-Zeeman SSOFs. Here, the condition $\langle\Phi(x)\rangle=0$ could be achieved by alternating the relative rotation angle of consecutive TMD layers, exploiting the anti-periodicity of the valley-Zeeman effect, $\lambda_{\textrm{vz}}(\theta)=-\lambda_{\textrm{vz}}(\theta\pm\pi/3)$ [43, 44]. The ensuing SSOF in this case strongly renormalizes the group velocity of wavepackets that propagate parallel to the SSOF direction. The leading correction to Eq. (3) is given by $\delta\varepsilon_{{\bf k}s}^{(n=\pm 1)}=\pm\mu\Lambda_{\textrm{vz}}\cos^{2}\theta_{\mathbf{k}}$, with $\Lambda_{\textrm{vz}}=\lambda_{\textrm{vz}}^{2}/(\hbar vG_{1})$, yielding an anisotropic energy dispersion and SB curvature. Interestingly, the valley-Zeeman SSOF leads to an overall decrease in the SB curvature magnitude, which is reflected in the spin Hall conductivity [Fig. (3)(c)]. This is also at odds with the expected behavior in (standard) proximitized graphene, where the spin-Hall conductivity has a non-monotonic behavior with $\lambda_{\text{vz}}$, with $\lambda_{\text{vz}}\neq 0$ [49] being essential to observe the SHE.
Figure 3: (a) Momentum-space distribution of the spin Berry curvature of the massless Dirac mini-bands ($n=\pm 1$). (b) Same as in (a) for the mini-bands $n=\pm 2$. SSOF parameters as in Fig. 2. (c) Room-temperature $\sigma^{z}_{xy}$ as a function of the chemical potential for selected twisting-induced SOC modulations: pure Rashba SSOFs with amplitudes as indicated (solid lines) and admixed Rashba-valley-Zeeman SSOFs with $\lambda_{\text{vz}}=\lambda_{\text{R}}/2$ (dashed line) and $\lambda_{\text{vz}}=\lambda_{\text{R}}$ (dot-dashed line).
@@
In closing, we have shown that the spatial patterning of symmetry-breaking spin-orbit fields gives rise to rich physics beyond that of conventional superlattices, most notably the emergence of 2D massless Dirac fermions with anomalous electrodynamic responses. The proposed periodic modulation of interface-induced SOC is within reach of current nano-fabrication methods, and is likely to have broad applications beyond those described in this work.
@@ I.1 Model and numerical approach
with $U(x)=u\Phi(x)$, $H_{\textrm{R}}(x)=\lambda_{R}\Phi(x)\,(\sigma_{x}\otimes s_{y}-\sigma_{y}\otimes s_{x})$, $H_{\textrm{vz,$\tau$}}(x)=\tau\lambda_{\textrm{vz}}\Phi(x)\,\sigma_{0}\otimes s_{z}$, and $H_{\textrm{KM}}(x)=\lambda_{\textrm{KM}}\Phi(x)\,\sigma_{z}\otimes s_{z}$. The wavefunctions in valley $\tau=\pm 1$ are 4-component spinors of the form $\Psi_{\tau}({\bf r})=(\psi_{\tau\uparrow}^{A}({\bf r}),\psi_{\tau\downarrow}^{A}({\bf r}),\psi_{\tau\uparrow}^{B}({\bf r}),\psi_{\tau\downarrow}^{B}({\bf r}))^{\textrm{t}}$. Moving to reciprocal space, the eigenproblem formally reduces to solving an infinite set of coupled equations for the plane-wave amplitudes $\{\psi_{\mathbf{k}s}^{\sigma}\}$:
$\hbar v|\mathbf{k}|e^{-i\sigma\theta_{\mathbf{k}}}\psi_{\mathbf{k}s}^{-\sigma}+\sum_{p\in\mathbb{Z}}\left[\left(u+s\sigma\lambda_{\textrm{KM}}+s\lambda_{\textrm{vz}}\right)\Phi_{G_{p}}\psi_{\mathbf{k}-\mathbf{G}_{p},s}^{\sigma}+i(s-\sigma)\lambda_{\textrm{R}}\Phi_{G_{p}}\psi_{\mathbf{k}-\mathbf{G}_{p},-s}^{-\sigma}\right]=E\psi_{\mathbf{k}s}^{\sigma}\,,$ (6)
@@
In our calculations, the summation over Fourier components in Eq. (6) is truncated to a finite, but large, number of terms ($|p|\leq N$). The resulting system of equations is solved numerically yielding $d=4(2N+1)$ bands ${\varepsilon_{n\mathbf{k}}}$ and associated 4-component eigenvectors ${\psi_{n\mathbf{k}}}$.
@@ I.2 Spatial profile of the periodic perturbation
where $\Phi$ is the amplitude and $R(x)=\Theta(x+\ell/2)\Theta(\ell/2-x)$. Here, $\Theta$ is the Heaviside step function, $a_{\text{S}}$ is the lattice width, $\ell$ is the barrier width ($\ell<a_{\text{S}}$), and $r=\ell/a_{\text{S}}$ (for a square wave $r=0.5$). For pure sinusoidal modulations, we use $\Phi(x)=\Phi_{c}+\Phi\cos(G_{1}x).$
@@ II.1 Numerics
  • spin-degenerate Dirac cones emerge at low energy (see remark below);
  • the Dirac cones are isotropic around $\bf k=0$ (see Fig. 6(a));
@@
The first point above is rigorously true for square-wave perturbations (i.e., $r=0.5$). For KP profiles with $r\neq 0.5$, the spin degeneracy is perturbatively lifted (see Eq. (11)). We also verified that sine-wave SSOFs behave similarly to square-wave SSOFs but with less pronounced features. This is easily explained by comparing their Fourier coefficients. For a square wave profile, $\Phi_{\pm G_{1}}^{\textrm{square}}=(2/\pi)\Phi$, while for sine waves: $\Phi_{\pm G_{1}}^{\textrm{sine}}=(1/2)\Phi$.
@@ II.2 Perturbation theory
$\varepsilon_{{\bf k}\sigma}^{(2)}\approx\sigma\hbar v\left[1-\mu\frac{\sin^{2}\theta_{{\bf k}}S^{2}+\cos^{2}\theta_{{\bf k}}Z^{2}+Y^{2}+\sin\theta_{{\bf k}}SY+2\sin\theta_{{\bf k}}\cos\theta_{{\bf k}}XY}{\hbar^{2}v^{2}G_{1}^{2}}\right]|{\bf k}|.\$ (10)
where $\mu$ is a geometric factor equal to $1$ for a sine-wave modulation and $\sum_{n=1}^{\infty}(4r/n){{}^{2}}{\rm sinc{{}^{2}}}(n\pi r)$ for a KP profiles with zero spatial average. We note that the gapless nature of the spectrum has its origin in the chiral phase factors $e^{i\theta_{{\bf k}}}$ of the unperturbed eigenstates as first discussed in Ref. [59] for pure scalar perturbations. Generalizing these results to the $4\times 4$ Dirac Hamiltonian in Eq. (5), containing the SSOFs, yields
@@ II.3 Third-order perturbation theory versus numerics
$\varepsilon_{{\bf k}s\sigma}^{(3)}\approx\varepsilon_{{\bf k}\sigma}^{(0)}-\mu\sigma\frac{\lambda_{\text{R}}^{2}}{\hbar vG_{1}^{2}}\left(|{\bf k}|\,-s\frac{\tilde{\mu}}{\mu}\frac{\lambda_{\text{R}}}{\hbar vG_{1}^{2}}\ |{\bf k}|^{2}\right),\,\,\,\,\,\,\,\,\,\,\,\Delta_{\text{spin}}(\mathbf{k})=2\tilde{\mu}|\lambda_{\textrm{R}}|\left(\frac{\lambda_{\text{R}}}{\hbar vG_{1}^{2}}\right)^{2}\ |{\bf k}|^{2},$ (12)
where $\tilde{\mu}$ is a geometric factor, $\tilde{\mu}=\sideset{}{{}_{n,m}^{\prime}}{\sum}4(2r)^{3}{\rm sinc}(n\pi r){\rm sinc}(m\pi r){\rm sinc}((n+m)\pi r)/[n^{2}(m+n)^{2}]$ (the primed summation is used to exclude the cases $n=-m$ and $n,m=0$). In accord with the numerics, we see that $\tilde{\mu}=0$ for $r=0.5$ (spin-degenerate states) and $\tilde{\mu}\neq 0$ for $r\neq 0.5$. The perturbative results reproduce well the numerics within their regime of validity, $|u|,|\lambda_{\text{KM}}|,|\lambda_{\text{R}}|,|\lambda_{\text{vz}}|\ll\pi\hbar v/a_{\text{S}}$ ($\approx 20$ meV for $a_{\text{S}}=100$ nm).
@@ II.4 Zero-energy solutions: analytical results
The spinor components can be cast as $\psi_{s}^{\sigma}({\bf r})=\phi_{s}^{\sigma}(x)e^{ik_{y}y}$, where $k_{y}$ is the wavevector component perpendicular to the superlattice direction. The Dirac equation for zero-energy states ($\hbar\equiv 1$) yields

### v2

@@ Abstract
The enhanced SB curvature of the SSOF-induced mini-bands indicates that the 2D van der Waals metamaterials proposed here support large spin Hall responses. To confirm this, we compute the intrinsic spin Hall conductivity ($\sigma_{ij}^{z}$) from the flux of SB curvature using standard methods [8] (see also Methods). According to linear response theory, the $z$-polarised spin current density generated by an external electric field is $\mathbf{j}_{s}^{z}=\sum_{i,j=x,y}\sigma_{ij}^{z}E_{j}\mathbf{e}_{j}$, where $E_{j}$ are the field components and $\textbf{e}_{i}$ is the unit vector along the $i$-axis. As shown in Fig. 3(c), the spin Hall response has a strong energy dependence and can reach sizeable values on the order of $e/4\pi$ for typical values of proximity-induced SOC at room temperature. This behaviour is robust to imperfections in the SSOF even when the vanishing global average condition, $\langle\Phi(x)\rangle=0$, is not exactly met. We verified this with different types of SOC and SSOF spatial patterns $\Phi(x)$. For example, the spin Hall conductivity presented in Fig. 3(c) is found to vary by less than 10% in the presence of a spatially uniform Rashba-like SOC as large as 50% of the SSOF magnitude itself; see SM for additional details. Moreover, at variance with 2D conductors subject to the usual uniform Rashba effect [77, 49], the spatial dependence of the Rashba SSOF protects our quantum geometry-driven spin Hall effect from exact cancellations due to impurity-scattering corrections. In fact, a semiclassical conservation law for expectation values involving the spin current can be derived in the vein of Ref. [49] yielding $\langle H_{\textrm{so}}(x)v_{i}^{z}\rangle=0$, with delimited-⟨$\langle$..$\rangle$ … ⟩ denoting a quantum and disorder average. For a uniform Rashba field, this relation (which holds in the presence of arbitrary non-magnetic impurity potentials) implies $\langle v_{i}^{z}\rangle=0$ in steady-state conditions and thus $\mathbf{j}_{s}^{z}=0$. However, in our system, the condition $\langle v_{i}^{z}\rangle=0$ is circumvented due to the oscillatory nature of $H_{\textrm{so}}(x)$. The SSOF-driven SHE thus appears to be more robust than its counterpart in standard Rashba-coupled graphene.
We now briefly address the case of 2D metamaterials with concurrent Rashba-type and valley-Zeeman SSOFs. Here, the condition $\langle\Phi(x)\rangle=0$ could be achieved by alternating the relative rotation angle of consecutive TMD layers, exploiting the anti-periodicity of the valley-Zeeman effect, $\lambda_{\textrm{vz}}(\theta)=-\lambda_{\textrm{vz}}(\theta\pm\pi/3)$ [43, 44]. The ensuing SSOF, in this case, strongly renormalises the group velocity of wavepackets that propagate parallel to the SSOF direction. The leading correction to Eq. (LABEL:eq:En3-perturbation-theory) is given by $\delta\varepsilon_{(n=\pm 1),{\bf k}s}=\pm\xi\Lambda_{\textrm{vz}}{\,\sin^{2}\theta_{\mathbf{k}}}$, with $\Lambda_{\textrm{vz}}=\lambda_{\textrm{vz}}^{2}/(\hbar v{G_{1}^{2})}$, yielding an anisotropic dispersion and SB curvature at low energies (see SM). The valley-Zeeman SSOF leads to an overall decrease in the SB curvature magnitude, which is reflected in the spin Hall conductivity [Fig. (3)(c)]. This is also at odds with the expected behaviour in (standard) proximitised graphene, where the spin Hall conductivity has a non-monotonic behaviour with $\lambda_{\text{vz}}$, with $\lambda_{\text{vz}}\neq 0$ [49, 50] being essential to observe the SHE.
@@
In closing, we have shown that the spatial patterning of symmetry-breaking spin-orbit fields gives rise to rich physics beyond that of conventional superlattices, in particular, the emergence of 2D massless Dirac fermions with anomalous electrodynamic responses. The proposed periodic modulation of interface-induced SOC is within reach of current nano-fabrication methods and is likely to have broad applications beyond those described in this work.
@@ I.1 Model and numerical approach
with $U(x)=u\Phi(x)$, $H_{\textrm{R}}(x)=\lambda_{R}\Phi(x)\,(\sigma_{x}\otimes s_{y}-\sigma_{y}\otimes s_{x})$, $H_{\textrm{vz,$\tau$}}(x)=\tau\lambda_{\textrm{vz}}\Phi(x)\,\sigma_{0}\otimes s_{z}$, and $H_{\textrm{KM}}(x)=\lambda_{\textrm{KM}}\Phi(x)\,\sigma_{z}\otimes s_{z}$ (see main text for definitions). The wavefunctions in valley $\tau=\pm 1$ are 4-component spinors of the form $\Psi_{\tau}({\bf r})=(\psi_{\tau\uparrow}^{A}({\bf r}),\psi_{\tau\downarrow}^{A}({\bf r}),\psi_{\tau\uparrow}^{B}({\bf r}),\psi_{\tau\downarrow}^{B}({\bf r}))^{\textrm{t}}$ for $\tau=1$ and $\Psi_{\tau}({\bf r})=(-\psi_{\tau\uparrow}^{B}({\bf r}),-\psi_{\tau\downarrow}^{B}({\bf r}),\psi_{\tau\uparrow}^{A}({\bf r}),\psi_{\tau\downarrow}^{A}({\bf r}))^{\textrm{t}}$ for $\tau=-1$. Moving to reciprocal space, the eigenproblem formally reduces to solving an infinite set of coupled equations for the plane-wave amplitudes $\{\psi_{\mathbf{k}s}^{\sigma}\}$ for each valley:
$\hbar v|\mathbf{k}|e^{-i\sigma\theta_{\mathbf{k}}}\psi_{\mathbf{k}s}^{-\sigma}+\sum_{p\in\mathbb{Z}}\left[\left(u+s\sigma\lambda_{\textrm{KM}}+s\tau\lambda_{\textrm{vz}}\right)\Phi_{G_{p}}\psi_{\mathbf{k}-\mathbf{G}_{p},s}^{\sigma}+i(s-\sigma)\lambda_{\textrm{R}}\Phi_{G_{p}}\psi_{\mathbf{k}-\mathbf{G}_{p},-s}^{-\sigma}\right]=E\psi_{\mathbf{k}s}^{\sigma}\,,$ (6)
@@
The summation over Fourier components in Eq. (6) is truncated to a finite, but large, number of terms ($|p|\leq N$). The resulting system of coupled equations is solved numerically yielding $d=4(2N+1)$ bands ${\varepsilon_{n\mathbf{k}}}$} and associated 4-component eigenvectors ${\psi_{n\mathbf{k}}}$}.
@@ I.2 Spatial profile of the periodic perturbation
where $\Phi$ is the SSOF amplitude and $R(x)=\Theta(x+\ell/2)\Theta(\ell/2-x)$. Here, $\Theta$ is the Heaviside step function, $a_{\text{S}}$ is the lattice width, $\ell$ is the barrier width ($\ell<a_{\text{S}}$), and $r=\ell/a_{\text{S}}$ (for a square wave $r=0.5$). For pure sinusoidal modulations, we use $\Phi(x)=\Phi_{c}+\Phi\cos(G_{1}x).$
@@ II.1 Numerics
  • spin-degenerate Dirac cones emerge at low energy (for $r=0.5$);
  • the Dirac cones are isotropic around $\bf k=0$ (see also Fig. 6(a));
@@
We also verified that sine-wave SSOFs behave similarly to square-wave SSOFs but with less pronounced features. This is easily explained by comparing their Fourier coefficients. For a square wave profile, $\Phi_{\pm G_{1}}^{\textrm{square}}=(2/\pi)\Phi$, while for sine waves: $\Phi_{\pm G_{1}}^{\textrm{sine}}=(1/2)\Phi$.
@@ II.2 Perturbation theory
$\varepsilon_{{\bf k}\sigma}^{(2)}$$\approx$ $\sigma\hbar v\left[1-\xi\frac{\sin^{2}\theta_{{\bf k}}S^{2}+\cos^{2}\theta_{{\bf k}}Z^{2}+Y^{2}+\sin\theta_{{\bf k}}SY+2\sin\theta_{{\bf k}}\cos\theta_{{\bf k}}XY}{\hbar{{}^{2}}v{{}^{2}}G_{1}{{}^{2}}}\right]|{\bf k}|.$ (10)
where $\xi$ is a geometric factor equal to $1$ for a sine-wave modulation and $\sum_{n=1}^{\infty}(4r/n){{}^{2}}{\rm sinc{{}^{2}}}(n\pi r)$ for a KP profiles with zero spatial average. We note that the gapless nature of the spectrum has its origin in the chiral phase factors $e^{i\theta_{{\bf k}}}$ of the unperturbed eigenstates as first discussed in Ref. [60] for pure scalar perturbations.
Generalizing these results to the $4\times 4$ Dirac Hamiltonian, containing the SSOFs, yields
@@
The group velocity for graphene under the combined action of a Kane-Mele SSOF and a periodic potential $U(x)$ of the main text can be directly derived from Eq. (11) by setting $\lambda_{\textrm{R}}=\lambda_{\textrm{vz}}=0$.
@@ II.3 Third-order perturbation theory versus numerics
$\varepsilon_{{\bf k}s\sigma}^{(3)}\approx\varepsilon_{{\bf k}\sigma}^{(0)}-\xi\sigma\frac{\lambda_{\text{R}}^{2}}{\hbar vG_{1}^{2}}\left(|{\bf k}|\,-s\frac{\tilde{\xi}}{\xi}\frac{\lambda_{\text{R}}}{\hbar vG_{1}^{2}}\ |{\bf k}|^{2}\right),\,\,\,\,\,\,\,\,\,\,\,\Delta_{\text{spin}}(\mathbf{k})=2\tilde{\xi}|\lambda_{\textrm{R}}|\left(\frac{\lambda_{\text{R}}}{\hbar vG_{1}^{2}}\right)^{2}\ |{\bf k}|^{2},$ (12)
where $\tilde{\xi}$ is a geometric factor, $\tilde{\xi}=\sideset{}{{}_{n,m}^{\prime}}{\sum}4(2r)^{3}{\rm sinc}(n\pi r){\rm sinc}(m\pi r){\rm sinc}((n+m)\pi r)/[n^{2}(m+n)^{2}]$ (the primed summation is used to exclude the cases $n=-m$ and $n,m=0$). In accord with the numerics, we see that $\tilde{\xi}=0$ for $r=0.5$ (spin-degenerate states) and $\tilde{\xi}\neq 0$ for $r\neq 0.5$. The perturbative results reproduce well the numerics within their regime of validity, $|u|,|\lambda_{\text{KM}}|,|\lambda_{\text{R}}|,|\lambda_{\text{vz}}|\ll\pi\hbar v/a_{\text{S}}$ ($\approx 20$ meV for $a_{\text{S}}=100$ nm).
@@ II.4 Zero-energy solutions: analytical results
Here, we provide explicit solutions for the $K$ valley (note that the expressions for $\tau=-1$ can be obtained from those below by simply inverting the sign of $\lambda_{\text{sv}}$). The spinor components can be cast as $\psi_{s}^{\sigma}({\bf r})=\phi_{s}^{\sigma}(x)e^{ik_{y}y}$, where $k_{y}$ is the wavevector component perpendicular to the superlattice direction. The Dirac equation for zero-energy states ($\hbar\equiv 1$) yields
@@
III Impact of spatially uniform components of SOC
Figure 8 shows the numerically calculated electronic structure of a graphene subject to a Rashba SSOF with non-zero global average, $\langle\Phi(x)\rangle\neq 0$. Compared to the case of a Rashba SSOF with $\langle\Phi(x)\rangle=0$ discussed in detail in the main text [see also panels (a) and (b) of Fig. 4 above, we see that the main modifications introduced by the spatially uniform Rashba component, $\lambda_{\textrm{Rc}}$, are twofold:
  • lifting of spin degeneracy
  • suppression of Dirac cones in favour of massive low-energy mini-bands.
(a) (b)
Refer to caption Refer to caption
Figure 8: Mini-band energy dispersion $\varepsilon(k_{x},0)$ (a) and $\varepsilon(0,k_{y})$ (b) of graphene subject to a square wave ($r=0.5$) Rashba modulation ($\lambda_{\text{R}}=20$ meV) with a uniform component corresponding to $\lambda_{\textrm{Rc}}=2$ meV; only positive energy branches are shown. Other parameters: $a_{\text{S}}=100$ nm. Bare energy dispersion is represented by dashed red lines as a guide to the eye.
The impact of a uniform Rashba component on the room-temperature spin-Hall conductivity is shown in Fig. 9 (a). We see that the spin-Hall response has a non-monotonic behavior with $\lambda_{\text{Rc}}$, with $|\sigma_{yx}^{z}|$ increasing slowly with $\lambda_{\text{Rc}}$ up to $\lambda_{\text{Rc}}\approx\lambda_{\text{R}}/2$ and then quickly decreasing beyond that. The response for a system subject to an admixed Rashba-valley-Zeeman SSOF containing a uniform valley-Zeeman component, $\lambda_{\textrm{vzc}}$, is shown in Fig. 9 (b). In this case, we also find that the spin-Hall response shows a non-monotonic behavior with $\lambda_{\text{vzc}}$, with $|\sigma_{yx}^{z}|$ changing slightly with $\lambda_{\text{vzc}}$ up to $\lambda_{\text{vzc}}\approx 2\lambda_{\text{\text{vz}}}$, while for values much greater it decreases significantly. These results show that $\sigma^{z}_{yx}$ is robust against fabrication imperfections leading to small uniform SOC components in the SSOF field.
(a) (b)
Refer to caption Refer to caption
Figure 9: Spin-Hall conductivity for square-wave ($r=0.5$) SSOFs with non-zero global average as a function of chemical potential for $k_{B}T=26$ meV and selected values of the uniform SOC components, $\lambda_{\textrm{Rc}}$ and $\lambda_{\textrm{zvc}}$, as indicated, for two scenarios: (a): Pure Rashba-type SSOF with $\lambda_{\text{R}}=20$ meV; and (b): Admixed Rashba–valley-Zeeman SSOF with $\lambda_{\text{R}}=20$ meV, $\lambda_{\textrm{Rc}}=10$ meV, and $\lambda_{\text{vz}}=10$ meV. Blue solid line of panel (a) and red dashed line of panel (b) correspond to blue solid line and blue dashed line of Fig. 3(c) in the main text, respectively. Other parameters as in Fig. 8.